OpenStudy (anonymous):

After a concert for 1,600 people, only 8 people said they would not go to see the band again. What percent of the people who went to the concert said they would not go to see the band again? 200 50% 8% 0.5%

OpenStudy (galacticplunder):

Well, what percent is 8 out of 1600?

OpenStudy (anonymous):

128? @GalacticPlunder

OpenStudy (mathstudent55):

Percent is a ratio out of 100. You are given a ratio of 8 out of 1600. You want to know what ratio out of 100 that equals. Set up a proportion: 8 out of 1600 equals x out of 100 \(\dfrac{8}{1600} = \dfrac{x}{100} \)

OpenStudy (mathstudent55):

Now solve the proportion for x by cross multiplying.

OpenStudy (anonymous):

0.5?

OpenStudy (galacticplunder):

Yes

OpenStudy (mathstudent55):

\(1600x = 8 \times 100\) \(1600x = 800\) \(x = \dfrac{800}{1600} \) \(x = 0.5\) Answer: 0.5% You are correct.